OpenStudy (anonymous):

Among all triangles of hypotenuse 10 units, determines the size of the area is maximized.

OpenStudy (australopithecus):

I haven't done this type of problem in a long time but I will tryto be helpful You have to take the derivative of A = ab/2 h^2 = a^2 + b^2 solve for one of the variables, h^(2) - b^(2) = a^2 (h^(2) - b^(2))^(1/2) = a \[ A(b) = \frac{(h^2 - b^2)^{1/2}b}{2}\] You can take the derivative, \[\frac{dA}{db} A(b) = \frac{dA}{db} \frac{(h^2 - b^2)^{1/2}b}{2}\] Find where, A'(b) = 0 and A'(b) is undefined but not undefined in the original function then make a table and see where the function is max and min, the value that gives you the minimum is equal to the value of b that gives the smallest area Do the same with variable a, then sub it into Pythagorean theorem and see if it gives you 10

OpenStudy (anonymous):

thank you very much :)